Speedy Auto Repairs uses a job-order costing system. The companys direct materials consist of replacement parts installed in customer vehicles, and its direct labor consists of the mechanics hourly wages. Speedys overhead costs include various items, such as the shop managers salary, depreciation of equipment, utilities, insurance, and magazine subscriptions and refreshments for the waiting room The company applies all of its overhead costs to jobs based on direct labor-hours. At the beginning of the year, t made the following estimates: Direct labor-hours required to support estimated output Fixed overhead cost Variable overhead cost per direct labor-hour 24,000 $288,000 $ 1.00 es Required: 1. Compute the predetermined overhead rate. 2. During the year, Mr. Wilkes brought in his vehicle to replace his brakes: spark plugs, and tires. The following information was available with respect to his job: Direct materials Direct labor cost Direct labor-hours used9 $ 615 $208 Compute Mr. Wilkes total job cost. selling prices using a markup percentage of 50% of its total job cost, then how much would it have 3. If Speedy establishes its Mr. Wilkes? Complete the question by entering your answers in the tabs given belovw Required 1 Required 2 Required 3
U3L, menn much would it have cha Complete the question by entering your answers in the tabs given below Required 1 Required 2 Required 3 Compute the predetermined overhead rate. (Round your answer to 2 decimal places.) rate per DLH Re Required 2>
Cumpute Mr. Wilkes total job cost. 31 Speedy establishes its selling prices using a markup percentage of 50% of its total ob cost, then how much would it have changed Mr. Wilkes? Complete the question by entering your answers in the tabs given belovw Required1Required 2 Required 3 Compute Mr. Wilkes, total Job cost. (Round your intermediate calculations to.2 decimal p aces Direct materials Direct labor Overhead applied Total cost assigned to Mr. Wikes S Required 1 Required 3
l, hen TnOw Huch would it have Complete the question by entering your answers in the tabs given below. Required 1 Required 2 Required 3 if Speedy establishes its selling prices using a markup percentage of 50% of its total job cost, then how much would it have Amount charged to Mr Wikes Required 2
Answer #1

1) Predetermine overhead rate = (288000/24000)+1 = $13 per direct labor hour

2) Calculate total product cost

Direct material 615
Direct labor 208
Overhead applied (9*13) 117
Total cost assigned to Mr.Wikes 940

3) Amount charged to Mr. Wikes = 940+(940*50%) = $1410